How much must you deposit each year into your retirement account starting now and continuing through year 10 if you want to be able to withdraw $100,000 per year forever, beginning 28 years from now? Assume the account earns interest at 12% per year.

The amount to be deposited is determined to be______ $  .
Answer #1

Ans:

the Annual worth of capital investment will be
A=P*i

The Interest rate is 12%
the Number of periods from now continuing till Year 10 is 11
Let Annual Deposit be A

A(p/a,12%,11) (F/P,12%,28)=100,000/.12
A(5.9377)(23.8839)=833,333.3333
a(141.8154)=833,333.3333
A=833,333.33/141.8154
A=5876.1836
The amount to be deposited is determined to be $__5876.1836